A 40% salt solution and a 85% salt solution are combined yielding 30L of 70% salt solution . What are the respective volumes of the 40% and 85% solutions?
Answer by lwsshak3(11628)

A 40% salt solution and a 85% salt solution are combined yielding 30L of 70% salt solution . What are the respective volumes of the 40% and 85% solutions?
***
let x=volume of 40% solution
30-x=volume of 85% solution
40%x+85%(30-x)=70%*30
.4x+25.5x-.85x=21
-.45x=4.5
x=10
30-x=20
volume of 40% solution=10 L
Volume of 85% solution=20 L